OpenStudy (anonymous):

A chemist needs 9 liters of 50% salt solution. All she has available is a 20% salt solution and a 70% salt solution. How much of each of the two solutions should she mix to obtain her desired solution? (Pleaseee do a step-by-step)

OpenStudy (anonymous):

x+y=9 .20x+.70y=4.5 2x+7y=45 2x+2y=18 5y=27 y=27/5 x=18/5 She needs 18/5 liters 20% solution and 27/5 liters 70% solution.

OpenStudy (anonymous):

let the chemist take x and y liters of each of the solution respectively, then x+y = 9------ equation I and , (20% of x) + ( 70% of y) = 50 % of 9 or 0.2x + 0.7 y = 4.5 or 2x + 7y = 45-------equation II Now, solve equation I and II to get x= 3.6 and y=5.4

OpenStudy (anonymous):

Thank you both soo much you were really a big help!